# SPECIAL RELATIVITY

### v9

@@ 16.3 XFEL Radiation Setup
The most elementary of the effect that represents a crucial test of the correct coupling fields and particles is a problem involves the production of coherent undulator radiation by modulated ultrarelativistic electron beam kicked by a weak dipole field before entering a downstream undulator. We want to study the process of emission of coherent undulator radiation from such setup.
The key element of a XFEL source is the udulator, which forces the electrons to move along curved periodical trajectories. There are two popular undulator configurations: helical and planar. To understand the basic principles of undulator source operation, let us consider the helical undulator (it is interesting to note that the first XFEL experiment demonstrating the apparent wavefront readjusting used helical undulator (3)).
@@
where $k_{w}=2\pi/\lambda_{w}$ is the undulator wavenumber and $\vec{e}_{x,y}$ are unit vectors directed along the $x$ and $y$ axes. We neglected the transverse variation of the magnetic field. It is necessary to mention that in XFEL engineering we deal with a very high quality of the undulator systems, which have a sufficiently wide good-field-region, so that our studies, which refer to a simple model of undulator field nevertheless yields a correct quantitative description in large variety of practical problems. The Lorentz force $\vec{F}=-e\vec{v}\times\vec{B}_{w}/c$ is used to derive the equation of motion of electrons with charge $-e$ and mass $m$ in the presence of magnetic field
@@
Let us consider a modulated ultrarelativistic electron beam moving alone the $z$ axis in the field of the helical undulator. In the present study we introduce the following assumptions. First, without kick the electrons move along constrained helical trajectories in parallel with the $z$ axis. Second, electron beam density at the undulator entrance is simply
@@
where $a=\mathrm{const.}$ In other words we consider the case in which there are no variation in amplitude and phase of the density modulation in the transverse plane. Under these assumptions the transverse current density may be written in the form
@@
Even through the measured quantities are real, it is generally more convenient to use complex representation, starting with real $\vec{j}_{\perp}$, one defines the complex transverse current density:
@@
Having defined the sources, we now should consider the electrodynamics problem. Maxwell equations can be manipulated mathematically in many ways in order to yield derived equations more suitable for certain applications. For example, from Maxwell equations Eq.(8) we can obtain an equation which depends only on the electric field vector $\vec{E}$ (in Gaussian units):
@@
Once the charge and current densities $\rho$ and $\vec{j}$ are specified as a function of time and position, this equation allows one to calculate the electric field $\vec{E}$ at each point of space and time. Thus, this nonhomogeneous wave equation is the complete and correct formula for radiation. However we want to apply it to still simpler circumstance in which second term (or, the current term) in the right-hand side provides the main contribution to the value of the radiation field. It is relevant to remember that our case of interest is the coherent undulator radiation and the divergence of this radiation is much smaller compared to the angle $1/\gamma$. It can be shown that when this condition is fulfilled the gradient term, $4\pi c^{2}\vec{\nabla}\rho$, in the right-hand side of the nonhomoheneous wave equation can be neglected. Thus we consider the wave equation
@@
We wish to examine the case when the phase velocity of the current wave is close to the velocity of light. This requirement may be met under resonance condition $\omega/c=\omega/v_{z}-k_{w}$. This is the condition for synchronism between the transverse electromagnetic wave and the fast transverse current wave with the propagation constant $\omega/v_{z}-k_{w}$. With the current wave traveling with the same phase speed as electromagnetic wave, we have the possibility of obtaining a spatial resonance between electromagnetic wave and electrons. If this the case, a cumulative interaction between modulated electron beam and transverse electromagnetic wave in empty space takes place. We are therefore justified in considering the contributions of all the waves except the synchronous one to be negligible as long as the undulator is made of a large number of periods.
Here follows an explanation of the resonance condition which is elementary in the sense that we can see what is happening physically. The field of electromagnetic wave has only transverse components, so the energy exchange between the electron and electromagnetic wave is due to transverse component of the electron velocity. For effective energy exchange between the electron and the wave, the scalar product $-e\vec{v}_{\perp}\cdot\vec{E}$ should be kept nearly constant along the whole undulator length. We see that required synchronism $k_{w}+\omega/c-\omega/v_{z}=0$ takes place when the wave advances the electron beam by the wavelength at one undulator period $\lambda_{w}/v_{z}=\lambda/(c-v_{z})$, where $\lambda=2\pi/\omega$ is the radiation wavelength. This tells us that the angle between the transverse velocity of the particle $\vec{v}_{\perp}$ and the vector of the electric field $\vec{E}$ remains nearly constant. Since $v_{z}\simeq c$ this resonance condition may be written as $\lambda\simeq\lambda_{w}/(2\gamma_{z}^{2})=\lambda_{w}(1+K^{2})/(2\gamma^{2})$.
We will use an adiabatic approximation that can be taken advantage of, in all practical situations involving XFELs, where the XFEL modulation wavelength is much shorter than the electron bunch length $\sigma_{b}$, i.e. $\sigma_{b}\omega/c\gg 1$. Since we are interested in coherent emission around the modulation wavelength the theory of coherent undulator radiation is naturally developed in the space-frequency domain. In fact, in this case one is usually interested into radiation properties at fixed modulation frequency.
@@
This is just a mathematical description of the concept of Huygens’ secondary sources and waves, and is of course well-known, but we still recalled how it follows directly from the Maxwell’s equations. We may consider the amplitude of the beam radiated by plane of oscillating electrons as a whole to be the resultant of radiated spherical waves. This is because Maxwell’s theory has no intrinsic anisotropy. The electrons lying on the plane of simultaneity gives rise to spherical radiated wavelets, and these combine according to Huygens’ principle to form what is effectively a radiated wave. If the plane of simultaneity is the $xy$-plane (i.e. beam modulation wavefront is perpendicular to the $z$- axis), then the Huygens’ construction shows that plane wavefronts will be emitted along the $z$-axis.
@@
In summary: according to Maxwell’s electrodynamics, coherent radiation is always emitted in the direction normal to the modulation wavefront. We already stressed that Maxwell’s equations are valid only in a Lorentz reference frame, i.e. when an inertial frame where the Einstein synchronization procedure is used to assign values to the time coordinates. Einstein’s time order should be applied and kept in consistent way in both dynamics and electrodynamics. Our previous description implies quite naturally that Maxwell’s equations in the lab frame are compatible only with covariant trajectories $\vec{x}_{cov}(t)$, calculated by using Lorentz coordinates and, therefore, including relativistic kinematics effects.
@@
Let us go back to the modulated electron beam, kicked transversely with respect to the direction of motion, that was discussed before. Conventional particle tracking shows that while the electron beam direction changes after the kick, the orientation of the modulation wavefront stays unvaried. In other words, the electron motion and the wavefront normal have different directions. Therefore, according to conventional coupling of fields and particles that we deem incorrect, the coherent undulator radiation in the kicked direction produced in a downstream undulator is expected to be dramatically suppressed as soon as the kick angle is larger than the divergence of the output coherent radiation.
@@
In order to estimate the loss in radiation efficiency in the kicked direction according to the conventional coupling of fields and particles, we make the assumption that the spatial profile of the modulation is close to that of the electron beam and has a Gaussian shape with standard deviation $\sigma$. A modulated electron beam in an undulator can be considered as a sequence of periodically spaced oscillators. The radiation produced by these oscillators always interferes coherently at zero angle with respect to the undulator axis. When all the oscillators are in phase there is, therefore, strong emission in the direction $\theta=0$. If we have a triangle with a small altitude $r\simeq\theta z$ and long base $z$, than the diagonal $s$ is longer than the base. The difference is $\Delta=s-z\simeq z\theta^{2}/2$. When $\Delta$ is equal to one wavelength, we get a minimum in the emission. This is because in this case the contributions of various oscillators are uniformly distributed in phase from 00 to $2\pi$. In the limit for a small size of the electron beam, $\sigma\to 0$, the interference will be constructive within an angle of about $\Delta\theta\backsimeq\sqrt{c/(\omega L_{w})}=1/(\sqrt{4\pi N_{w}}\gamma_{z})\ll 1/\gamma$, where $L_{w}=\lambda_{w}N_{w}$ is the undulator length. In the limit for a large size of the electron beam, the angle of coherence is about $\Delta\theta\backsimeq c/(\omega\sigma)$ instead. The boundary between these two asymptotes is for sizes of about $\sigma_{\mathrm{dif}}\backsimeq\sqrt{cL_{w}/\omega}$. The parameter $\omega\sigma^{2}/(cL_{w})$ can be referred to as the electron beam Fresnel number. It is worth noting that, for XFELs, the transverse size of electron beam $\sigma$ is typically much larger than $\sigma_{\mathrm{dif}}$ (i.e electron beam Fresnel number is large). Thus, we can conclude that the angular distribution of the radiation power in the far zone has a Gaussian shape with standard deviation $\sigma_{\mathrm{\theta}}\backsimeq c/(\sqrt{2}\omega\sigma)$. However, still according to the conventional treatment, after the electron beam is kicked we have the already-mentioned discrepancy between direction of the electron motion and wavefront normal. Then, the radiation intensity along the new direction of the electron beam can be approximated as $I\backsimeq I_{0}\exp[-\theta_{k}^{2}/(2\sigma_{\mathrm{\theta}}^{2})]$, where $I_{0}$ is the on-axis intensity without kick and $\theta_{k}$ is the kick angle. The exponential suppression factor is due to the tilt of the modulation wavefront with respect to the direction of motion of the electrons.
@@
We presented a study of very idealized situation for illustrating the difference between conventional and covariant coupling of fields and particles. We solved the dynamics problem of the motion of a relativistic electrons in the prescribed force field of weak kicker magnet by working only up to the order of $\gamma\theta_{k}$. This approximation is of particular theoretical interest because it is relatively simple and at the same time forms the basis for understanding relativistic kinematic effects such as relativity of simultaneity.
@@
Let us discuss the region of validity of our small kick angle approximation $\theta_{k}\gamma\ll 1$. Since in XFELs the Fresnel number is rather large, we can always consider a kick angle which is relatively large compared to the divergence of the output coherent radiation, and, at the same time, it is relatively small compared to the angle $1/\gamma$. In fact, from $\omega\sigma^{2}/(cL_{w})\gg 1$, with some rearranging, we obtain $\sigma_{\theta}^{2}\simeq c^{2}/(\omega^{2}\sigma^{2})\ll c/(\omega L_{w})$. Then we recall that $\sqrt{c/(\omega L_{w})}=1/(\sqrt{4\pi N_{w}}\gamma_{z})\ll 1/\gamma$. Therefore, the first order approximation used to investigate the kicker setup in this chapter is of practical interest in XFEL engineering.
@@
It is one of the aims of this chapter is to demonstrate the kind of experimental predictions we are expecting from our corrected radiation theory. We worked out a very simple case in order to illustrate all the essential physical principles very clearly. Surprisingly, the first order approximation used to investigate the kicker setup in this section has also important practical applications.
@@
Above we have shown that our correct coupling of fields and particles predicts an effect in complete contrast to the conventional treatment. Namely, in the ultrarelativistic limit, the plane of simultaneity, that is wavefront orientation of the modulation, is always perpendicular to the electron beam velocity. As a result, we predict strong emission of coherent undulator radiation from the modulated electron beam in the kicked direction, Fig. 73.
@@
Figure 73: Basic setup for the experimental test of XFEL theory. The correct coupling fields and particles predicts an effect in complete contrast to the conventional treatment. According to the covariant approach, in the ultrarelativistic limit, the wavefront of modulation is always perpendicular to the electron beam velocity. As a result, the Maxwell’s equations predict strong emission of coherent undulator radiation from the modulated electron beam in the kicked direction. Experiments show that this prediction is, in fact, true.
@@
XFEL experts actually witnessed an apparent wavefront readjusting due to the relativistic kinematics effect, but they never drew this conclusion. In this book, we are actually first in considering the idea that results of the conventional theory of radiation by relativistically moving charges are not consistent with the special relativity. In previous literature, identification of the trajectories in the source part of the usual Maxwell’s equations with the trajectories calculated by conventional particle tracking in the (”single”) lab frame has always been considered obvious.
@@
In existing literature theoretical analysis is presented, of an XFEL driven by an electron beam with wavefront tilt, and this analysis is based on the exploitation of usual Maxwell’s equations and standard simulation codes. Using only a kicker setup (i.e. without undulator radiation setup) we can demonstrate that the coupling fields and particles in the conventional XFEL theory is intrinsically incorrect.
@@
The existing XFEL theory based on the use of the absolute time convention (i.e. old kinematics) for particle dynamics. Here we will give a simple proof of the conflict between conventional particle tracking and Maxwell’s electrodynamics. The purpose is to show how one can demonstrate in a simple way that the conventional XFEL theory is absolutely incapable of correctly describing the distribution of the electromagnetic fields from a fast moving modulated electron beam downstream the kicker.
@@
Under the Maxwell’s electrodynamics, the fields of a modulated electron beam moving with a constant velocity exhibit an interesting behavior when the velocity of charges approaches that of light: namely, in the space-time domain they resemble more and more close of a laser beam (see Appendix IV). In fact, for a rapidly moving modulated electron beam we have nearly equal transverse and mutually perpendicular electric and magnetic fields: in the limit $v\to c$ they become indistinguishable from the fields of a laser beam, and according to Maxwell’s equations, the wavefront of the laser beam is always perpendicular to the propagation direction. This is indeed the case for virtual laser-like radiation beam in the region upstream the kicker.
@@
Let us now consider the effect of the kick on the electron modulation wavefront. If we rely on the conventional particle tracking, the kick results in a difference between the directions of electron motion and the normal to the modulation wavefront, i.e. in a tilt of the modulation wavefront.
@@
This is already a conflict result, because we now conclude that, according to the conventional ”single frame” approach, the direction of propagation after the kick is not perpendicular to the radiation beam wavefront. In other words, the radiation beam motion and the radiation wavefront normal have different directions. The virtual radiation beam (which is indistinguishable from a real radiation beam in the ultrarelativistic asymptote) propagates in the kicked direction with a wavefront tilt. This is what we would get for the case when our analysis is based on the conventional particle tracking, and is obviously absurd from the viewpoint of Maxwell’s electrodynamics.
@@
In existing literature theoretical analysis is presented, of an XFEL driven by an electron beam with wavefront tilt, and this analysis is based on the exploitation of usual Maxwell’s equations and standard simulation codes. Using only a kicker setup (i.e. without undulator radiation setup) we demonstrated that the coupling fields and particles in the conventional XFEL theory is intrinsically incorrect.
@@
In existing XFEL theory the wavefront tilt is considered as real. However, there is a common mistake made in accelerator physics connected with the wavefront tilt. In the ultrarelativistic domain the wavefront tilt has no exact objective meaning. The angle of wavefront tilt depends on the choice of a procedure for clock synchronization in the lab frame, as a result of which it can be given any preassigned values within the interval $(0,\theta_{k})$. For instance, in the ultrarelativistic domain, the orientation of the modulation wavefront is always perpendicular to the electron beam velocity (i.e. $\theta_{tilt}=0$) when the evolution of the modulated electron beam is treated using Lorentz coordinates. No physical effects may depends on an arbitrary constant or an arbitrary function (5).
16.5 Final Remarks

### v10

@@ 20.3 XFEL Radiation Setup
The most elementary of the effects that represent a crucial test of the correct coupling fields and particles is a problem involving the production of coherent undulator radiation by modulated ultrarelativistic electron beam kicked by a weak dipole field before entering a downstream undulator. We want to study the process of emission of coherent undulator radiation from such a setup.
The key element of an XFEL source is the undulator, which forces the electrons to move along curved periodical trajectories. There are two popular undulator configurations: helical and planar. To understand the basic principles of undulator source operation, let us consider the helical undulator.
@@
where $k_{w}=2\pi/\lambda_{w}$ is the undulator wavenumber and $\vec{e}_{x,y}$ are unit vectors directed along the $x$ and $y$ axes. We neglected the transverse variation of the magnetic field. It is necessary to mention that in XFEL engineering we deal with very high-quality undulator systems, which have a sufficiently wide good-field-region, so that our studies, which refer to a simple model of undulator field nevertheless yield a correct quantitative description in a large variety of practical problems. The Lorentz force $\vec{F}=-e\vec{v}\times\vec{B}_{w}/c$ is used to derive the equation of motion of electrons with charge $-e$ and mass $m$ in the presence of magnetic field
@@
Let us consider a modulated ultrarelativistic electron beam moving alone on the $z$ axis in the field of the helical undulator. In the present study, we introduce the following assumptions. First, without kick the electrons move along constrained helical trajectories in parallel with the $z$ axis. Second, electron beam density at the undulator entrance is simply
@@
where $a=\mathrm{const.}$ In other words we consider the case in which there are no variations in amplitude and phase of the density modulation in the transverse plane. Under these assumptions, the transverse current density may be written in the form
@@
Even though the measured quantities are real, it is generally more convenient to use complex representation, starting with real $\vec{j}_{\perp}$, one defines the complex transverse current density:
@@
Having defined the sources, we now should consider the electrodynamics problem. Maxwell equations can be manipulated mathematically in many ways in order to yield derived equations more suitable for certain applications. For example, from Maxwell equations Eq.(8) we can obtain an equation that depends only on the electric field vector $\vec{E}$ (in Gaussian units):
@@
Once the charge and current densities $\rho$ and $\vec{j}$ are specified as a function of time and position, this equation allows one to calculate the electric field $\vec{E}$ at each point of space and time. Thus, this nonhomogeneous wave equation is the complete and correct formula for radiation. However, we want to apply it to still simpler circumstances in which the second term (or, the current term) on the right-hand side provides the main contribution to the value of the radiation field. It is relevant to remember that our case of interest is the coherent undulator radiation and the divergence of this radiation is much smaller compared to the angle $1/\gamma$. It can be shown that when this condition is fulfilled the gradient term, $4\pi c^{2}\vec{\nabla}\rho$, in the right-hand side of the nonhomogeneous wave equation can be neglected. Thus we consider the wave equation
@@
We wish to examine the case when the phase velocity of the current wave is close to the velocity of light. This requirement may be met under resonance condition $\omega/c=\omega/v_{z}-k_{w}$. This is the condition for synchronism between the transverse electromagnetic wave and the fast transverse current wave with the propagation constant $\omega/v_{z}-k_{w}$. With the current wave traveling at the same phase speed as the electromagnetic wave, we have the possibility of obtaining a spatial resonance between the electromagnetic wave and electrons. If this is the case, a cumulative interaction between the modulated electron beam and a transverse electromagnetic wave in empty space takes place. We are therefore justified in considering the contributions of all the waves except the synchronous one to be negligible as long as the undulator is made of a large number of periods.
Here follows an explanation of the resonance condition which is elementary in the sense that we can see what is happening physically. The field of electromagnetic wave has only transverse components, so the energy exchange between the electron and electromagnetic wave is due to a transverse component of the electron velocity. For effective energy exchange between the electron and the wave, the scalar product $-e\vec{v}_{\perp}\cdot\vec{E}$ should be kept nearly constant along the whole undulator length. We see that required synchronism $k_{w}+\omega/c-\omega/v_{z}=0$ takes place when the wave advances the electron beam by the wavelength at one undulator period $\lambda_{w}/v_{z}=\lambda/(c-v_{z})$, where $\lambda=2\pi/\omega$ is the radiation wavelength. This tells us that the angle between the transverse velocity of the particle $\vec{v}_{\perp}$ and the vector of the electric field $\vec{E}$ remains nearly constant. Since $v_{z}\simeq c$ this resonance condition may be written as $\lambda\simeq\lambda_{w}/(2\gamma_{z}^{2})=\lambda_{w}(1+K^{2})/(2\gamma^{2})$.
We will use an adiabatic approximation that can be taken advantage of, in all practical situations involving XFELs, where the XFEL modulation wavelength is much shorter than the electron bunch length $\sigma_{b}$, i.e. $\sigma_{b}\omega/c\gg 1$. Since we are interested in coherent emission around the modulation wavelength the theory of coherent undulator radiation is naturally developed in the space-frequency domain. In fact, in this case, one is usually interested in radiation properties at fixed modulation frequency.
@@
This is just a mathematical description of the concept of Huygens’ secondary sources and waves and is of course well-known, but we still recall how it follows directly from Maxwell’s equations. We may consider the amplitude of the beam radiated by the plane of oscillating electrons as a whole to be the resultant of radiated spherical waves. This is because Maxwell’s theory has no intrinsic anisotropy. The electrons lying on the plane of simultaneity give rise to spherical radiated wavelets, and these combine according to Huygens’ principle to form what is effectively a radiated wave. If the plane of simultaneity is the $xy$-plane (i.e. beam modulation wavefront is perpendicular to the $z$- axis), then Huygens’ construction shows that plane wavefronts will be emitted along the $z$-axis.
@@
In summary: according to Maxwell’s electrodynamics, coherent radiation is always emitted in the direction normal to the modulation wavefront. We already stressed that Maxwell’s equations are valid only in a Lorentz reference frame, i.e. when an inertial frame where the Einstein synchronization procedure is used to assign values to the time coordinates. Einstein’s time order should be applied and kept in a consistent way in both dynamics and electrodynamics. Our previous description implies quite naturally that Maxwell’s equations in the lab frame are compatible only with covariant trajectories $\vec{x}_{cov}(t)$, calculated by using Lorentz coordinates and, therefore, including relativistic kinematics effects.
@@
Let us go back to the modulated electron beam, kicked transversely with respect to the direction of motion, that was discussed before. Conventional particle tracking shows that while the electron beam direction changes after the kick, the orientation of the modulation wavefront stays unvaried. In other words, the electron motion and the wavefront normal have different directions. Therefore, according to the conventional coupling of fields and particles that we deem incorrect, the coherent undulator radiation in the kicked direction produced in a downstream undulator is expected to be dramatically suppressed as soon as the kick angle is larger than the divergence of the output coherent radiation.
@@
To estimate the loss in radiation efficiency in the kicked direction according to the conventional coupling of fields and particles, we make the assumption that the spatial profile of the modulation is close to that of the electron beam and has a Gaussian shape with standard deviation $\sigma$. A modulated electron beam in an undulator can be considered as a sequence of periodically spaced oscillators. The radiation produced by these oscillators always interferes coherently at zero angle with respect to the undulator axis. When all the oscillators are in phase there is, therefore, strong emission in the direction $\theta=0$. If we have a triangle with a small altitude $r\simeq\theta z$ and long base $z$, then the diagonal $s$ is longer than the base. The difference is $\Delta=s-z\simeq z\theta^{2}/2$. When $\Delta$ is equal to one wavelength, we get a minimum in the emission. This is because in this case the contributions of various oscillators are uniformly distributed in phase from 00 to $2\pi$. In the limit for small size of the electron beam, $\sigma\to 0$, the interference will be constructive within an angle of about $\Delta\theta\backsimeq\sqrt{c/(\omega L_{w})}=1/(\sqrt{4\pi N_{w}}\gamma_{z})\ll 1/\gamma$, where $L_{w}=\lambda_{w}N_{w}$ is the undulator length. In the limit for the large size of the electron beam, the angle of coherence is about $\Delta\theta\backsimeq c/(\omega\sigma)$ instead. The boundary between these two asymptotes is for sizes of about $\sigma_{\mathrm{dif}}\backsimeq\sqrt{cL_{w}/\omega}$. The parameter $\omega\sigma^{2}/(cL_{w})$ can be referred to as the electron beam Fresnel number. It is worth noting that, for XFELs, the transverse size of electron beam $\sigma$ is typically much larger than $\sigma_{\mathrm{dif}}$ (i.e. electron beam Fresnel number is large). Thus, we can conclude that the angular distribution of the radiation power in the far zone has a Gaussian shape with standard deviation $\sigma_{\mathrm{\theta}}\backsimeq c/(\sqrt{2}\omega\sigma)$. However, still according to the conventional treatment, after the electron beam is kicked we have the already-mentioned discrepancy between the direction of the electron motion and wavefront normal. Then, the radiation intensity along the new direction of the electron beam can be approximated as $I\backsimeq I_{0}\exp[-\theta_{k}^{2}/(2\sigma_{\mathrm{\theta}}^{2})]$, where $I_{0}$ is the on-axis intensity without kick and $\theta_{k}$ is the kick angle. The exponential suppression factor is due to the tilt of the modulation wavefront with respect to the direction of motion of the electrons.
@@
We presented a study of the very idealized situation to illustrate the difference between the conventional and covariant coupling of fields and particles. We solved the dynamics problem of the motion of relativistic electrons in the prescribed force field of a weak kicker magnet by working only up to the order of $\gamma\theta_{k}$. This approximation is of particular theoretical interest because it is relatively simple and at the same time forms the basis for understanding relativistic kinematic effects such as relativity of simultaneity.
@@
Let us discuss the region of validity of our small kick angle approximation $\theta_{k}\gamma\ll 1$. Since in XFELs, the Fresnel number is rather large, we can always consider a kick angle that is relatively large compared to the divergence of the output coherent radiation, and, at the same time, it is relatively small compared to the angle $1/\gamma$. In fact, from $\omega\sigma^{2}/(cL_{w})\gg 1$, with some rearranging, we obtain $\sigma_{\theta}^{2}\simeq c^{2}/(\omega^{2}\sigma^{2})\ll c/(\omega L_{w})$. Then we recall that $\sqrt{c/(\omega L_{w})}=1/(\sqrt{4\pi N_{w}}\gamma_{z})\ll 1/\gamma$. Therefore, the first-order approximation used to investigate the kicker setup in this chapter is of practical interest in XFEL engineering.
@@
It is one of the aims of this chapter is to demonstrate the kind of experimental predictions we are expecting from our corrected radiation theory. We worked out a very simple case in order to illustrate all the essential physical principles very clearly. Surprisingly, the first-order approximation used to investigate the kicker setup in this section has also important practical applications.
@@
Above we have shown that our correct coupling of fields and particles predicts an effect in complete contrast to the conventional treatment. Namely, in the ultrarelativistic limit, the plane of simultaneity, that is wavefront orientation of the modulation, is always perpendicular to the electron beam velocity. As a result, we predict a strong emission of coherent undulator radiation from the modulated electron beam in the kicked direction, Fig. 62.
@@
Figure 62: Basic setup for the experimental test of XFEL theory. According to the covariant approach, in the ultrarelativistic limit, the wavefront of modulation is always perpendicular to the electron beam velocity. As a result, Maxwell’s equations predict a strong emission of coherent undulator radiation from the modulated electron beam in the kicked direction. Experiments show that this prediction is, in fact, true.
@@
XFEL experts actually witnessed an apparent wavefront readjusting due to the relativistic kinematics effect, but they never drew this conclusion. In this book, we are actually first in considering the idea that the results of the conventional theory of radiation by relativistically moving charges are not consistent with special relativity. In previous literature, identification of the trajectories in the source part of the usual Maxwell’s equations with the trajectories calculated by conventional particle tracking in the (”single”) lab frame has always been considered obvious.
@@
In existing literature, theoretical analysis is presented, of an XFEL driven by an electron beam with wavefront tilt, and this analysis is based on the exploitation of usual Maxwell’s equations and standard simulation codes. Using only a kicker setup (i.e. without undulator radiation setup) we can demonstrate that the coupling fields and particles in the conventional XFEL theory are intrinsically incorrect.
@@
The existing XFEL theory is based on the use of the absolute time convention (i.e. old kinematics) for particle dynamics. Here we will give simple proof of the conflict between conventional particle tracking and Maxwell’s electrodynamics. The purpose is to show how one can demonstrate in a simple way that the conventional XFEL theory is incapable of correctly describing the distribution of the electromagnetic fields from a fast-moving modulated electron beam downstream the kicker.
@@
Under Maxwell’s electrodynamics, the fields of a modulated electron beam moving with a constant velocity exhibit an interesting behavior when the velocity of charges approaches that of light: namely, in the space-time domain they resemble more and more closer to a laser beam (see Appendix IV). In fact, for a rapidly moving modulated electron beam we have nearly equal transverse and mutually perpendicular electric and magnetic fields: in the limit $v\to c$ they become indistinguishable from the fields of a laser beam, and according to Maxwell’s equations, the wavefront of the laser beam is always perpendicular to the propagation direction. This is indeed the case for virtual laser-like radiation beam in the region upstream the kicker.
@@
Let us now consider the effect of the kick on the electron modulation wavefront. If we rely on conventional particle tracking, the kick results in a difference between the directions of electron motion and the normal to the modulation wavefront, i.e. in a tilt of the modulation wavefront.
@@
This is already a conflict result because we now conclude that, according to the conventional ”single frame” approach, the direction of propagation after the kick is not perpendicular to the radiation beam wavefront. In other words, the radiation beam motion and the radiation wavefront normal have different directions. The virtual radiation beam (which is indistinguishable from a real radiation beam in the ultrarelativistic asymptote) propagates in the kicked direction with a wavefront tilt. This is what we would get for the case when our analysis is based on conventional particle tracking, and is absurd from the viewpoint of Maxwell’s electrodynamics.
@@
In existing literature, theoretical analysis is presented, of an XFEL driven by an electron beam with wavefront tilt, and this analysis is based on the exploitation of usual Maxwell’s equations and standard simulation codes. Using only a kicker setup (i.e. without undulator radiation setup) we demonstrated that the coupling fields and particles in the conventional XFEL theory are intrinsically incorrect.
@@
20.5 Discussion